Update #9: Magnetometer has been restored and data is flowing. Other instruments status is TBD
Update #7: GOES-17 recovery is in progress. The anomaly has been addressed. Engineers are working towards a full recovery. Instrument data is expected to be restored over night.
Update #6: G17 recovery operations is still ongoing.
Update #5: G17 recovery operations still ongoing.
Update #4: G17 recovery operations still ongoing.
Update #3: Engineers have started G17 recovery operations.
Update #2: Engineers are onsite at NSOF and aggressively troubleshooting the problem. All GOES-17 ABI, MAG, SEISS, SUVI AND EXIS/EUV products are out due to this ongoing Anomaly.
Update #1: Engineers are onsite at NSOF to diagnose the GOES-17 Anomaly.
Topic: GOES-17 ABI L1b All Bands, All Products and derived products delivered to AWIPS, GRB, PDA
Date/Time Issued: July23, 2021 0527Z
Product or Data Impacted: GOES-17 ABI L1b All Bands, All Products and derived products All Scenes delivered to AWIPS, GRB, PDA.
Date/Time of Initial Impact: July 22, 2021 0537Z
Date/Time of Expected End: TBD
Length of the Outage/Event: TBD
Details/Specifics of Change: GOES-17 ABI L1b All Bands, All Products and derived products All Scenes delivered to AWIPS, GRB, PDA. OUTAGE due to G17 entering SAFEHOLD MODE, engineering investigating.
Contact information for further information: ESPC Operations at ESPCOperations@noaa.gov
or 301-817-3880 24x7 or firstname.lastname@example.org
during normal business hours
This message was sent by ESPC.Notification@noaa.gov
. You have been sent this and other notifications because you have opted in to receive it. If for any reason, you wish to unsubscribe, please contact ESPC Help Desk at ESPCOperations@noaa.gov
or (301) 817-3880. Please note: it may take up to two business days to process your unsubscribe request.